if you invested 30000 dollars in two simple-interest bank accounts and one payed 6% and the other payed 8% annual interest and after one year you had 220$, how much did you invest in each account?
Answer by mananth(16946)

I think there is a typo
either 30,000 and 2200
or3000 & 220
Part I 6.00% per annum ------------- Amount invested =x
Part II 8.00% per annum ------------ Amount invested = y
30000
Interest----- 2200.00

Part I 6.00% per annum ---x
Part II 8.00% per annum ---y
Total investment
x + 1 y= 30000 -------------1
Interest on both investments
6.00% x + 8.00% y= 2200
Multiply by 100
6 x + 8 y= 220000.00 --------2
Multiply (1) by -6
we get
-6 x -6 y= -180000.00
Add this to (2)
0 x 2 y= 40000
divide by 2
y = 20000
Part I 6.00% $ 10000
Part II 8.00% $ 20000

CHECK
10000 --------- 6.00% ------- 600.00
20000 ------------- 8.00% ------- 1600.00
Total -------------------- 2200.00

m.ananth@hotmail.ca